# Knife-Edge Diffraction of Scalar and Vector Vortex Light

### v1

@@ I Introduction
Light provides a canvas that may be structured in its intensity, phase, and polarization among other degrees of freedom. Since the beginning of this millennium, the arrival of spatial light modulators, digital mirror displays and other specialized devices in optical laboratories has allowed us to generate and analyze a wide range of optical structures. Optical vortices, be it phase vortices that carry optical angular momentum (OAM) within homogeneously polarized light, or vector vortices, featuring polarization singularities, capture the attention of researchers more than other optical structures.
@@
Phase vortices are characterized by a helical phase structure that twists around the optical axis, resulting in a phase singularity (or dislocation) in the center of the beam [1, 2, 3, 4, 5, 6], they have been the subject of numerous investigations, summarized in various reviews [7, 8, 9, 10, 11].
@@
Vector light, characterized by inhomogeneous polarization profiles, is a more recent addition to the structured light family. Vector light may feature vector vortices with topological singularities in the polarization field [12]. Typically these arise when superimposing different phase vortices in orthogonal polarization components. For details on their characterization as well as their significance for light matter interaction and optical information encoding we refer the reader to recent reviews [13, 14, 15, 16].
@@
While the interference of phase structures results in intensity modulations, the interference of polarization structures results in polarization modulations across the beam profile. The presence of optical vortices is therefore visible in any experimental situation that involves interference in its widest sense, including e.g. diffraction or scattering. This needs to be considered when interpreting measurements, but in turn may also be utilized for the detection of vortex light.
@@
Direct interference with its mirror image or a reference beam was perhaps the first method identified to measure OAM [17, 18] – and the characteristic spiraling or fork-like intensity patterns are still widely used [19]. Even single plano-concave lenses can be used as rudimentary interferometric devices by superposing the reflections from the front and back surfaces of the lens [20]. Mode-conversion may be understood as a less obvious form of (intra-beam) interference, and can also be used to analyze vortex light, including astigmatic systems (such as tilted or cylindrical lenses) [21, 22], as well as more sophisticated devices that work on the single photon level [23, 24]. Diffraction offers an alternative route, including multiple-pinhole diffraction [25], single or double-slit diffraction [26, 27, 28], and diffraction through triangular or circular apertures [29, 30, 31], as well as different types of gratings [32, 33, 34, 35], single-point detection [36], spatial light modulators [37], and specially designed nanoholes [38].
@@
Knife-edge diffraction can be understood as the interference of two superimposing waves: the geometrical wave from the input beam and the boundary diffraction wave from the diffracting object. It has been beautifully demonstrated that this effect can be used to analyze the phase structure of a diffractive object [39]. In reverse, diffraction off a known object (i.e. the knife-edge) can reveal the light beam’s phase structure. Two decades ago, diffraction of a phase vortex by a knife-edge has been used to probe the handedness of its topological charge [40]. Like other single-beam interferometers, knife-edge diffraction is self-referencing, and hence provides some stability to aberrations and external vibrations.
@@ II.2 Simulation
We model the propagation and diffraction of Eq. (1) based on ideal LG modes. The real-space grid is defined based on the experimental parameters of the imaging system, and the corresponding reciprocal space grid is made sufficiently large to allow propagation in the Fourier domain.
The knife-edge is modeled as a binary mask, where regions of the beam beyond a specified position (defined in terms of the beam waist $w_{0}$) are blocked. The beam is also confined by a circular aperture to simulate the finite size of optical elements.
@@
The simulation results are visualized using intensity, phase and polarization maps and compared with experimental results, with details of the various visualizations and analytical tools given in the following section III.
@@ III.1 Experimental setup
$\vec{u}(r,\phi)=\frac{1}{\sqrt{2}}\left(e^{i\alpha/2}{\rm LG}_{0}^{m}\hat{h}+e^{-i\alpha/2}{\rm LG}_{0}^{-m}\hat{v}\right),$ (5)
@@ III.2 Knife-edge diffraction of phase vortex beams
Figure 2: Observed diffraction pattern for light with $m=+1$ (left image) and $m=-1$ (right image) as a function of the lateral position of the knife edge at positions (a)$-2w_{0}$, (b) $-w_{0}$, (c) $-w_{0}/2$, (d) 0, (e) $w_{0}/4$, (f) $w_{0}$, with respect to the beam’s center. The circular insets show the corresponding simulations for $m=+1$, with the intensity profile displayed in gray levels, the phase in hue colors, and the dotted lines showing the edge position (those for $m=-1$ are horizontal mirror images).
Before the knife-edge intersects the beam significantly, the $m=\pm 1$ phase vortices are intact as shown in Fig. 2(a), but as the knife intersects the light (b-c), the characteristic fork pattern emerges. The diffraction patterns for ±mplus-or-$\pm m$ are horizontal mirror images of each other, and the sign of the vortex can clearly be identified from the fork orientation, confirming a previous study [40]. For $m=+1$ ($m=-1$) more light invades the geometric shadow in the lower (upper) half plane, in agreement with the clockwise (anticlockwise) tilt of the Poynting vectors with skew angles ±$\pm 1/(kr)$. Once the knife edge passes the vortex position (e,f), the fork disappears again, however the tilt of the diffraction pattern remains.
@@
We analyze the presence and disappearance of the point vortex in more detail in Fig. 3. This figure displays the accumulated phase along a closed path around the beam center, i.e. the singularity position of the input beam, divided by $2\pi$. This value measures the vortex charge, corresponding to the beam’s OAM.
As expected, the phase initially exhibits a $2\pi$ variation around the vortex (for $m=1$). As the edge covers the vortex, the accumulated phase goes to zero, corresponding to the disappearance of the fork-like fringes in Fig. 2. Analogous behavior occurs also for negative and higher-order vortices (not shown).
@@
We have also measured and simulated knife-edge diffraction for higher order phase vortices, as illustrated in Fig. 4, choosing a lateral knife-edge position of $-w_{0}/2$ which gives the clearest fork patterns. We observe $m$-pronged forks, demonstrating that knife-edge diffraction is a simple and effective method to determine the value and sign of the topological charge of a vortex beam. Our measurements also indicate that the extent to which the diffracted beam penetrates the geometric shadow region is influenced by $|m|$. Again, we may understand this intuitively by considering the tilt of the Poynting vector in the presence of a vortex, which is proportional to $|m|$. This phenomenon can be interpreted as a redistribution of energy due to the vortex topology, which influences the diffraction pattern’s behavior [48, 49, 50].
@@ IV Conclusion
Knife-edge diffraction does not receive the same attention as other diffraction phenomena like the double slit or gratings, despite its honorable mention in undergraduate optics courses and its importance for wireless communication. Here we have presented a detailed analysis of knife-edge diffraction for scalar and vector vortices. For phase vortices, we observe fork dislocations within the familiar Fresnel diffraction fringes that depend on both the sign and magnitude of the vortex charge. Our observations agree with simulations based on the ASM. For vector vortices, we find a modulation of the ellipticity close to the shadow region. The effects in both scalar and vector vortices illustrate the energy flow given by the tilted Poynting vector of the phase vortex beams.
While our experiment was optimized to demonstrate the fork dislocations in the knife-edge diffraction pattern as cleanly as possible, a qualitative observation of the diffraction pattern can be achieved very quickly by inserting a blade partly into a laser beam, offering a simple and effective method to determine the charge and sign of a vortex. The method should be transferrable to any beams that can scatter, offering potential vortex measurements for media where other detection methods are costly or not yet available.
@@ Acknowledgements.
R.M.A. received funding via Fraunhofer CAP, award 322761-01. This study was financed in part by the Coordenação de Aperfeiçoamento de Pessoal de Nível Superior - Brasil (CAPES) - Finance Code 001.

### v2

@@ I Introduction
Light provides a canvas that may be structured in its intensity, phase, and polarization among other degrees of freedom. Since the beginning of this millennium, the arrival of spatial light modulators, digital mirror displays and other specialized devices in optical laboratories has allowed us to generate and analyze a wide range of optical structures. Optical vortices, be it phase vortices carrying optical angular momentum (OAM) within homogeneously polarized light, or vector vortices featuring polarization singularities, capture the attention of researchers more than other optical structures.
@@
Phase vortices are characterized by a helical phase structure that twists around the optical axis, resulting in a phase singularity (or dislocation) in the center of the beam [1, 2, 3, 4, 5, 6]. They have been the subject of numerous investigations, summarized in various reviews [7, 8, 9, 10, 11].
@@
Vector light, characterized by inhomogeneous polarization profiles, is a more recent addition to the structured light family. Vector light may feature vector vortices with topological singularities in the polarization field [12]. Typically, these arise when superimposing different phase vortices in orthogonal polarization components. For details on their characterization as well as their significance for light matter interaction and optical information encoding we refer the reader to recent reviews [13, 14, 15, 16].
@@
While the interference of phase structures results in intensity modulations, the interference of polarization structures results in polarization modulations across the beam profile. The presence of optical vortices is therefore visible in any experimental situation that involves interference in its widest sense, including e.g. diffraction or scattering. This needs to be considered when interpreting measurements, and in turn may be utilized for the detection of vortex light.
@@
Direct interference with its mirror image or a reference beam was perhaps the first method identified to measure the OAM of a light beam [17, 18] – and the characteristic spiraling or fork-like intensity patterns are still widely used [19]. Even single plano-concave lenses can serve as rudimentary interferometric devices by superposing the reflections from the front and back surfaces of the lens [20]. Mode-conversion may be understood as a less obvious form of (intra-beam) interference, and its use for the analysis of vortex light includes astigmatic systems (such as tilted or cylindrical lenses) [21, 22], as well as more sophisticated devices that work on the single photon level [23, 24]. Diffraction offers an alternative route, including multiple-pinhole diffraction [25], single or double-slit diffraction [26, 27, 28, 29], and diffraction through triangular or circular apertures [30, 31, 32], as well as different types of gratings [33, 34, 35, 36], single-point detection [37], spatial light modulators [38], and specially designed nanoholes [39]. In this paper we will elaborate on knife-edge diffraction as an efficient and experimentally simple topological method to identify the charge and sign of scalar vortices, and furthermore to illustrate features of vector vortices.
@@
Knife-edge diffraction can be understood as the interference of two superimposing waves: the geometrical wave from the input beam and the boundary diffraction wave from the diffracting object. It has been beautifully demonstrated that this effect can be used to analyze the phase structure of a diffractive object [40]. In reverse, diffraction off a known object (i.e. the knife-edge) can reveal critical information about the light beam’s phase structure [41]. Two decades ago, diffraction of a phase vortex by a knife-edge has been used to probe the handedness of its topological charge [42]. Like other single-beam interferometers, knife-edge diffraction is self-referencing, and hence provides some stability to aberrations and external vibrations. The sign and charge of an optical phase vortex is revealed by the orientation of the fork dislocation present in the diffraction pattern, making knife-edge diffraction an experimentally simple topological method for the analysis of optical vortices.
@@ II.2 Simulation
We model the propagation and diffraction of Eq. (1) based on ideal LG modes. The real-space grid is defined based on the experimental parameters of the imaging system, and the corresponding reciprocal space grid is made sufficiently large to allow propagation in the Fourier domain. The knife-edge is modeled as a binary mask, where regions of the beam beyond a specified position (defined in terms of the beam waist $w_{0}$) are blocked. The beam is also confined by a circular aperture to simulate the finite size of optical elements.
@@
The simulation results for intensity, phase and polarization patterns offer a numerical validation of the experimental measurements, serving as a qualitative confirmation of the interpretation of this phenomenon through the various visualizations and analytical tools given in the following section III. While the intensity patterns reveal the spatial distribution of optical power, phase maps highlight the structure of vortices, and polarization maps provide insight into the interplay between spatial truncation and polarization states. These visualizations are key to understanding the impact of truncation on the beam’s propagation and diffraction properties.
@@ III.1 Experimental setup
$\mathbf{u}(r,\phi)=\frac{1}{\sqrt{2}}\left(e^{i\alpha/2}{\rm LG}_{0}^{m}\hat{h}+e^{-i\alpha/2}{\rm LG}_{0}^{-m}\hat{v}\right),$ (5)
@@ III.2 Knife-edge diffraction of phase vortex beams
Figure 2: Observed diffraction pattern for light with $m=+1$ (left image) and $m=-1$ (right image) as a function of the lateral position of the knife edge at positions (a)$-2w_{0}$, (b) $-w_{0}$, (c) $-w_{0}/2$, (d) 0, (e) $w_{0}/4$, (f) $w_{0}$, with respect to the beam’s center. The circular insets show the corresponding simulations for $m=+1$, with the intensity profile displayed in gray levels, the phase in hue colors, and the dotted lines showing the edge position (those for $m=-1$ are horizontal mirror images).To guide the eye, we have indicated the “fork” with gray lines in (b).
Before the knife-edge intersects the beam significantly, the $m=\pm 1$ phase vortices are intact as shown in Fig. 2(a), but as the knife intersects the light (b-c), the characteristic fork pattern emerges. The diffraction patterns for ±mplus-or-$\pm m$ are horizontal mirror images of each other, featuring a fork dislocation with $|m|+1=2$ prongs. The sign of the vortex can clearly be identified from the fork orientation, confirming a previous study [42]. For $m=+1$ ($m=-1$) more light invades the geometric shadow in the lower (upper) half plane, in agreement with the clockwise (anticlockwise) energy circulation[50] as determined by the tilt of the Poynting vectors with skew angles ±$\pm 1/(kr)$. Once the knife edge passes the vortex position (e,f), the fork disappears again, however the tilt of the diffraction pattern remains.
@@
We analyze the presence and disappearance of the point vortex in more detail in Fig. 3. This figure displays the accumulated phase along a closed path around the beam center, i.e. the singularity position of the input beam, divided by $2\pi$. This value measures the vortex charge, corresponding to the OAM of the input beam  [^2] We note that, as the propagation direction changes during diffraction, the OAM or the input light does not coincide with that of the diffracted beam - for small diffraction angles considered here it is however a good approximation.. As expected, the phase initially exhibits a $2\pi$ variation around the vortex (for $m=1$). As the edge covers the vortex, the accumulated phase goes to zero, corresponding to the disappearance of the fork-like fringes in Fig. 2. Analogous behavior occurs also for negative and higher-order vortices (not shown).
@@
We have also measured and simulated knife-edge diffraction for higher order phase vortices, as illustrated in Fig. 4, choosing a lateral knife-edge position of $-w_{0}/2$ which gives the clearest fork patterns. For an input beam with topological charge $m$, we observe $|m|+1$-pronged forks. Together with our results from Fig. 2, this demonstrates that knife-edge diffraction is a simple and effective method to determine the value and sign of the topological charge of a vortex beam from the shape and orientation of the fork dislocation. Our measurements also indicate that the extent to which the diffracted beam penetrates the geometric shadow region is influenced by $|m|$. Again, we may understand this intuitively by considering the tilt of the Poynting vector in the presence of a vortex, which is proportional to $|m|$. This phenomenon can be interpreted as a redistribution of energy due to the vortex topology, which influences the diffraction pattern’s behavior [52, 53, 54, 50].
@@ IV Conclusion
Knife-edge diffraction does not receive the same attention as other diffraction phenomena like the double slit or gratings, despite its honorable mention in undergraduate optics courses and its impact on wireless communication. Here we have presented a detailed analysis of knife-edge diffraction for scalar and vector vortices. For phase vortices, we observe fork dislocations within the familiar Fresnel diffraction fringes that depend on both the sign and magnitude of the vortex charge. Our observations agree with simulations based on the ASM. For vector vortices, we find a modulation of the ellipticity close to the shadow region. The effects in both scalar and vector vortices illustrate the energy flow given by the tilted Poynting vector of the phase vortex beams.
While our experiment was optimized to demonstrate the fork dislocations in the knife-edge diffraction pattern as cleanly as possible, a qualitative observation of the diffraction pattern can be achieved very quickly by inserting a blade partly into a laser beam, offering a simple and effective method to determine the charge and sign of a vortex. Counting dislocations of a fork singularity is practical only for small OAM numbers for two reasons: Automating the process of analyzing the fringe pattern, while in principle possible, is a non-trivial task. The clarity of fringes deteriorates in the presence of noise, especially for higher order OAM beams. We further note that the method is easily transferrable to the single-photon regime or for beams with non-integer OAM. Knife-edge diffraction, as a method of identifying topological charge, has been demonstrated for electron vortices [55],, and more generally, the method should be transferrable to any beams that can scatter, offering potential vortex measurements for media where other detection methods are costly or not yet available.
@@ Acknowledgements.
R.M.A. received funding via Fraunhofer CAP, award 322761-01. SF-A acknowledges support through the QuantERA II Programme, with funding received via the EU H2020 research and innovation programme under Grant No. 101017733 and assoicated support from EPSRC under Grant No. EP/Z000513/1 (V-MAG). A.K.F. received funding via CAPES, Process number 88887.837237/2023-00. This study was financed in part by the Coordenação de Aperfeiçoamento de Pessoal de Nível Superior - Brasil (CAPES) - Finance Code 001.